I am not having a good week where the math problems are concerned. Here it is;
Solve for S
1/R=1/S+1/T
I am very lost, help...

Answer by Alan3354(69443)

Solve for S
1/R=1/S+1/T
--------------
1/s = 1/r - 1/t
1/s = (t-r)/rt
s = rt/(t-r)